OpenStudy (anonymous):

Solve the proportion for the given variable. t/14 = 49/35

OpenStudy (bahrom7893):

t = 49*14/35

OpenStudy (bahrom7893):

t = 49*2/5

OpenStudy (nenadmatematika):

t=(14*49)/35

OpenStudy (bahrom7893):

t = 19.6

OpenStudy (anonymous):

t=98.5

OpenStudy (anonymous):

bahrom7893 i had 19.6 to but it was was wrong

OpenStudy (bahrom7893):

That's correct, then the question is wrong..

OpenStudy (nenadmatematika):

it's correct chinagirl check if you wrote the question correctly :D

OpenStudy (anonymous):

\[686/35 = 19.6, what form do they want the answer \in?\]

OpenStudy (anonymous):

What format do you want the answer in?

OpenStudy (anonymous):

19 and 21/35. all are equivlent

OpenStudy (anonymous):

t = 49 * 14/ 35 = 19.6